# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ 5 Discussion of Convergence Rate and Acceleration
$P_{ij}^{*}=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm if}\leavevmode\nobreak\ \Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}}.\end{aligned}\right.$
Here, we finish the proof. ∎
Remark 9.
Under abovementioned discussion, $g_{2}(x)$ is strongly convex with parameter $\mu_{g_{2}}=\min\{a_{1},a_{2}\}$. It is not hard to show that $g_{\mathcal{Q}}(P)=\sum_{i,j}w_{ij}g_{2}(P_{ij})$ is strongly convex with parameter $\mu_{g_{\mathcal{Q}}}=\min_{i,j}w_{ij}\mu_{g_{2}}$, if for all $i,j$, $w_{ij}\neq 0$.
@@
1:Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},y_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{k}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}>0,\gamma_{0}>0$, smoothness parameter $L_{f}$
2:Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
3:for $k=0,1,2,...$ do
4:     Compute $\alpha_{k}$ by (44)
5:     Compute $\eta_{g,k},\tilde{y}_{k},\tilde{\eta}_{f,k},\tilde{v}_{k}$ by (14), (15), (16), (17)
6:     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),\leavevmode\nobreak\ d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
7:     Update $v_{k+1}\leftarrow$Inner Layer Solver
8:     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
9:     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
10:     Update $\widetilde{P}_{k+1}$ by (43)
11:     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
12:     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (18)
13:     if Stopping Criterion==absent$=$ =True then
14:         return $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
15:     end if
16:end for
Interestingly, the only difference between Algorithm 2 and Algorithm 3 is the update of $\widetilde{P}_{k+1}$. However, the former holds $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate, while the latter holds $\mathcal{O}\left(\frac{1}{k^{2}}\right)$ convergence rate. Note that Nesterov’s accelerated gradient descent is an extension of vanilla gradient descent by introducing an additional sequence and modifying the fundamental framework; Algorithm 2 and Algorithm 3 have a completely consistent framework. Specifically, the following theorem is proposed.
Theorem 6.
Provided that for all $i,j$, $w_{ij}\neq 0$, and $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, under the initial condition
$\gamma_{0}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}=\mu_{g_{\mathcal{Q}}}>0$
and the condition
@@
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
@@
where $\mu_{g_{\mathcal{Q}}}=\min_{i,j}w_{ij}\mu_{g_{2}}$, provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
Proof.
Noticing that $f_{1}(\widetilde{W})$ is a linear function, $\nabla f_{1}$ is Lipschitz continuous with Lipschitz parameter $L_{f_{1}}=0$. In addition, $g_{\mathcal{Q}}$ is $\mu_{g_{\mathcal{Q}}}$-strongly convex. Then, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
@@ 6 Directly Optimizing $\ell_{0}$-penalty
In the above sections, our primary objective is to minimize the $\ell_{0}$ norm (See Definition 9) of $W_{2}$; and indeed, the $\ell_{1}$-penalty and the piecewise quadratic penalty function can be considered as a convex relaxation of the $\ell_{0}$ norm (see Figure 1). In this section, the $\ell_{0}$ sparsity promoting function is considered directly, which makes the proposed optimization problem nonconvex and discontinuous. We will show that a series of coordinatewise convex (but nonconvex) problems can serve as an apporximation for the nonconvex problem at hand. Then, the BSUM framework is employed, allowing for the effective handling of the coordinatewise convex problem. Finally, some theorems are proposed to illustrate the variational properties of $\ell_{0}$-penalty induced problem. We first introduce some standard definitions below.
@@
$z\in{\rm dom}f\subseteq\mathbb{R}^{m}$ is the coordinatewise minimizer of $f$ with respective to the coordinates in $\mathbb{R}^{m_{1}},\dots,\mathbb{R}^{m_{n}}$, $m_{1}+\dots+m_{n}=m$ if
$f(z+d_{k}^{0})\geq f(z)\leavevmode\nobreak\ \forall d_{k}\in\mathbb{R}^{m_{k}}\leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ z+d_{k}^{0}\in{\rm dom}f\leavevmode\nobreak\ \forall k=1,\dots,n,$
@@
The subsequent analysis in this section is predicated upon the following assumption.
@@
Definition 9.
For $x=(x_{1},\dots,,x_{n})^{\top}\in\mathbb{R}^{n}$, consider the $\ell_{0}$-norm function $f:\mathbb{R}^{n}\to\mathbb{R}$ given by
$f(x)=\|x\|_{0}\equiv\#\{i\colon x_{i}\neq 0\}.$
That is, $\|x\|_{0}$ is the number of nonzero elements in $x$.
Remark 10.
In reality, $\ell_{0}$-norm is not a norm, since it does not satisfy the homogeneity property. Note that $\ell_{0}$-norm is a discontinuous and nonconvex function, but can be viewed as a sum of closed functions
$\|x\|_{0}=\sum_{i=1}^{n}I(x_{i}),$
where $I:\mathbb{R}\to\{0,1\}$ is given by
$I(y)=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ y=0,\\ &1,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{otherwise}.\end{aligned}\right.$
Hence, $\ell_{0}$-norm is closed. Whatever, significant challenges arise when $\ell_{0}$-penalty is introduced into the optimization problems.
@@
$\min_{\widetilde{W}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}$ (45)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,\dots,M.$
Denote
@@
$\min_{\widetilde{W}}\leavevmode\nobreak\ \langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}.$ (P)
@@
$\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|y|)=I(y)=\left\{\begin{aligned} &0,\leavevmode\nobreak\ \leavevmode\nobreak\ y=0,\\ &1,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{otherwise}.\end{aligned}\right.$
@@
$\min_{\widetilde{W}}\leavevmode\nobreak\ \langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right).$ ($\textbf{P}_{\sigma}$)
@@
Proof.
By Property $\mathcal{D}$, $g_{\sigma}(x)$ is a proper closed and convex function. Hence, by Theorem 4.3 of [1], $g^{*}_{\sigma}(y)$ $(y<0)$ is a closed convex function. By Theorem 4.8 of [1], we have $g_{\sigma}^{**}=g_{\sigma}$. Hence, $g_{\sigma}=\sup\limits_{y\leq 0}\{yx-g_{\sigma}^{*}(y)\}$. Equivalently, $f_{\sigma}(x)=\inf\limits_{y\leq 0}\{-yx+g^{*}_{\sigma}(y)\}$. The proof is complete. ∎
@@
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}\leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})$ ($\textbf{P}^{\prime}_{\sigma}$)
$+\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\gamma g_{\sigma}^{*}(-y).$
@@
$r_{1}(\widetilde{W})=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W}),$
@@
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}\leavevmode\nobreak\ H_{\sigma}(\widetilde{W},y)=r_{1}(\widetilde{W})+r_{2,\sigma}(y)+h(\widetilde{W},y).$ ($\textbf{P}^{\prime\prime}_{\sigma}$)
It is important to acknowledge that $r_{1},r_{2}$ are closed convex functions, but $h$ is nonconvex and nonsmooth. It seems that, compared to optimization problem ($\textbf{P}_{\sigma}$), ($\textbf{P}^{\prime\prime}_{\sigma}$) remains a nonconvex problem but introduces auxiliary variable $y$, which even makes ($\textbf{P}^{\prime\prime}_{\sigma}$) more challenging to solve. Interestingly, it is evident that $H_{\sigma}(\widetilde{W},y)$ is closed convex with regard to $\widetilde{W}$ and $y$, respectively. This observation suggests that by employing block coordinate descent, ($\textbf{P}^{\prime\prime}_{\sigma}$) can be converted to successive convex optimization problems. Specifically, given $y_{k-1}$, update $\widetilde{W}_{k}$; then given $\widetilde{W}_{k}$, update $y_{k}$, i.e.,
$\widetilde{W}_{k}\in\underset{\widetilde{W}\in\mathbb{R}^{p^{2}}}{\operatorname*{argmin}}\leavevmode\nobreak\ r_{1}(\widetilde{W})+h(\widetilde{W},y_{k-1}),$ ($\textbf{BP}\sigma-1$)
$y_{k}\in\underset{y\in\mathbb{R}_{+}^{mn}}{\operatorname*{argmin}}\leavevmode\nobreak\ r_{2,\sigma}(y)+h(\widetilde{W}_{k},y).$ ($\textbf{BP}\sigma-2$)
It is visible that both ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) may be classified as convex optimization problems. Our objective is then to solve ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternatively in order to approximate the global minimizers of nonconvex problem ($\textbf{P}^{\prime\prime}_{\sigma}$). Before discussing the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and ($\textbf{BP}\sigma-1$) ($\textbf{BP}\sigma-2$), we first propose the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and (P). Denote
@@
Proof.
Since for $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$
@@
for $W\in\mathbb{S}_{+}^{p}$. According to Exercise 7.32 of [1], the sequence $\{H_{\sigma_{i}}\}_{i\in\mathbb{N}}$ is eventually level-bounded. By noticing that $H_{\sigma_{i}}$ and $H$ are lower semi-continuous and proper, we finish the proof by Theorem 7.33 of [1]. ∎
Hence, theoretically, one can obtain the minimizers of optimization (P) by successively solving optimization problem ($\textbf{P}_{\sigma}$) (or, equivalently, ($\textbf{P}^{\prime\prime}_{\sigma}$)) with $\sigma\downarrow 0$. As mentioned above, we hope to solve optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) by block coordinate descent, i.e, finding the global optimal of convex optimization problems ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternately. Regretfully, since unseparable term $h(\widetilde{W},y)$ is nonconvex and nonsmooth, direct block coordinate descent method may cause stable cyclic behavior, whose gradient is bounded away from zero in the limiting path [39]. Hence, we will then modify the optimization problems ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$). Before that, we first introduce the auxiliary functions.
Definition 11.
  1. 1.
    $u_{i}(y_{i},y)=f(y)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall y\in\mathcal{X},\forall i$,
  2. 2.
    $u_{i}(x_{i},y)\geq f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i,$
  3. 3.
    $u_{i}^{\prime}(x_{i},y;d_{i})\big{|}_{x_{i}=y_{i}}=f^{\prime}(y;d)\leavevmode\nobreak\ \leavevmode\nobreak\ \forall d=(0,\dots,d_{i},\dots,0)\text{s.t. }y_{i}+d_{i}\in\mathcal{X}_{i}\forall i$,
  4. 4.
    $u_{i}(x_{i},y)$ is continuous in $(x_{i},y)\forall i$.
Remark 11.
For a continuous coordinatewise convex function $f$, auxiliary functions always exist. A trivial choice of auxiliary functions is
$u_{i}(x_{i},y)=f(y_{1},\dots,y_{i-1},x_{i},y_{i+1},\dots,y_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall x_{i}\in\mathcal{X}_{i},\forall y\in\mathcal{X},\forall i.$
Noticing that a convex function is differentiable in any direction at any interior point of its domain [36], the conditions 1-4 of the above definition naturally hold.
In [40], an unified algorithm framework is introduced to solve the following optimization problem
$\min f(x),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t. }x\in\mathcal{X},$ (48)
where $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{m}$, and $\mathcal{X}\subseteq\mathbb{R}^{m}$ is the Cartesian product of $n$ closed convex sets: $\mathcal{X}=\mathcal{X}_{1}\times\dots\times\mathcal{X}_{n}$ with $\mathcal{X}_{i}\in\mathbb{R}^{m_{i}}$ and $\sum_{i}m_{i}=m$.
  Algorithm 4 BSUM
1:Given : Feasible initial point $x^{0}\in\mathcal{X}$
2:Result: $x^{r}$
3:Set $r=0$
4:for $k=0,1,2,...$ do
5:     Let $r=r+1,i=(r\mod n)+1$
6:     Let $\mathcal{X}^{r}=\underset{x_{i}\in\mathcal{X}_{i}}{\operatorname*{argmin}}u_{i}(x_{i},x^{r-1})$
7:     Set $x^{r}_{i}$ to be an arbitrary element in $\mathcal{X}^{r}$
8:     Set $x_{k}^{r}=x_{k}^{r-1}\leavevmode\nobreak\ \forall k\neq i$
9:     if Stopping Criterion==absent$=$ =True then
10:         return $x^{r}$
11:     end if
12:end for
The above pseudocode of BSUM (Block Successive Upper-bound Minimization) framework introduced in [40] gives methodology for solving the optimization problem (48), where $\{u_{i}\}_{i=1}^{n}$ are auxiliary functions of $f$. It should be noted that the straight implementation of the BSUM framework is not viable due to the lack of knowledge on the auxiliary functions and the presence of multiple subproblems (see line 6 in Algorithm 4) that need to be addressed.
Specifically, back to our topic, optimization problem ($\textbf{BP}\sigma-2$) at hand remains unaltered. In other word, the auxiliary function $u_{2,\sigma}(y,X_{k})$ is selected as
$u_{2,\sigma}(y,X_{k})=\langle{\rm vec}(R),\widetilde{W}_{k}\rangle+r_{2,\sigma}(y)+h(\widetilde{W}_{k},y),$ (49)
where $X_{k}=(\widetilde{W}_{k},y_{k})$. ($\textbf{BP}\sigma-2$) is equivalent to
@@
Proof.
Optimality condition implies that
@@
The proof is complete. ∎
We then discuss the selection of $u_{1}(\widetilde{W},X_{k})$. The proximal term is introduced, i.e.,
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$ (52)

### v3

@@ 5 Discussion of Convergence Rate and Acceleration
$\hat{P}_{ij;k+1}^{*}=\left\{\begin{aligned} &0,~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}>\frac{\gamma w_{ij}b_{1}}{\rho_{k}},\\ &\frac{\rho_{k}\Omega_{ij}^{k}-\gamma w_{ij}b_{1}}{\gamma w_{ij}a_{1}+\rho_{k}},~{}~{}~{}~{}{\rm if}~{}\Omega_{ij}^{k}\leq\frac{\gamma w_{ij}b_{1}}{\rho_{k}}.\end{aligned}\right.$
[…]
@@
  Given : Initial point $\widetilde{W}_{0}\in\Gamma^{p},v_{0}\in\Gamma^{p},y_{0}\in\mathbb{R}^{mn},w_{0}\in\mathbb{R}^{mn},\lambda_{k}\in\mathbb{R}^{mn},\theta_{0}=1,\beta_{0}=\mu_{g_{\mathcal{Q}}},\gamma_{0}>0$, smoothness parameter $L_{f}$
  Result : $\widetilde{W}_{k}$, $\widetilde{P}_{k}$
  for $k=0,1,2,...$ do
     Compute $\alpha_{k}$ by (53)
     Compute $\eta_{g,k},\widetilde{y}_{k},\widetilde{\eta}_{f,k},\widetilde{v}_{k}$ by (22), (23), (24), (25)
     Determine $u_{k}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k}),~{}d_{k}=\nabla f_{1}(u_{k})+\mathcal{A}^{\top}\lambda_{k}$
     Update $v_{k+1}\leftarrow$Inner Layer Solver
     Update $\widetilde{W}_{k+1}=(1+\alpha_{k})^{-1}(\widetilde{W}_{k}+\alpha_{k}v_{k+1})$
     Update $\bar{\lambda}_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k})$
     Update $\widetilde{P}_{k+1}$ by (51)
     Update $\lambda_{k+1}=\lambda_{k}+\alpha_{k}/\theta_{k}(\mathcal{A}v_{k+1}+\mathcal{B}w_{k+1})$
     Update $\theta_{k+1},\gamma_{k+1},\beta_{k+1}$ by (26)
     if Stopping Criterion==absent$=$ =True then
        return  $\widetilde{W}_{k+1},\widetilde{P}_{k+1}$
     end if
  end for
Interestingly, the only difference between Algorithm 3 and Algorithm 4 is the update of $\widetilde{P}_{k+1}$. However, the former holds $\mathcal{O}\left(\frac{1}{k}\right)$ convergence rate, while the latter holds $\mathcal{O}\left(\frac{1}{k^{2}}\right)$ convergence rate. Note that Nesterov GD is an extension of vanilla GD by introducing an additional sequence and modifying the fundamental framework; Algorithm 3 and Algorithm 4 have a completely consistent framework. We have established a direct relationship between the acceleration of solving decentralized LQ problems and the design of a strongly convex sparsity promoting function, and to our best knowledge, this phenomenon has not been reported in the field of sparse optimization.
Corollary 3.
Provided that for all $i,j$, $w_{ij}\neq 0$, and $a_{1}>0,a_{2}>0,b_{1}<0<b_{2}$, under the initial condition $\gamma_{0}>0$, $\beta_{0}=\mu_{g_{\mathcal{Q}}}>0$ and the condition
@@
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},$
$\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
@@
provided that $\beta_{0}\leq\left\|\mathcal{B}\right\|^{2}$.
[…]
@@ 6 Directly Optimizing $\ell_{0}$-penalty
Reviewing the fundamental topic of this paper, we want to penalize the number of communication links among decentralized controllers by introducing the sparsity promoting function into the optimal control problem. In this section, the optimization problem
$\min_{W\in\mathbb{S}^{p}}{}{}$ $\langle R,W\rangle+\gamma\|V_{1}WV_{2}^{\top}\|_{0}$ (54)
${\rm s.t.}{}{}$ $W\in\mathscr{C}$
is directly studied, which is nonconvex and discontinuous. It is worth to highlight that penalizing the number of communication links among decentralized controllers is equivalent to the penalty term $\gamma\|W_{2}\|_{0}$ of problem (54), and the penalty promoting terms $\gamma g(W_{2})$ and $\gamma g_{\mathcal{Q}}(W_{2})$ introduced in the previous sections can be regarded as the convex relaxations of $\gamma\|W_{2}\|_{0}$. By the following discussions, we show that a series of coordinatewise convex problems can serve as an approximation for problem (54). Additionally, the variational properties of problem (54) are of concern. We first introduce some standard definitions below.
@@
$z\in{\rm dom}f\subseteq\mathbb{R}^{m}$ is the coordinatewise minimizer of $f$ with respective to the coordinates in $\mathbb{R}^{m_{1}},\dots,\mathbb{R}^{m_{n}}$, $m_{1}+\dots+m_{n}=m$ if $\forall k=1,\dots,n$
$f(z+d_{k}^{0})\geq f(z)~{}\forall d_{k}\in\mathbb{R}^{m_{k}}~{}{\rm with}~{}z+d_{k}^{0}\in{\rm dom}f,$
@@
The subsequent analysis in this section is based on the following assumption.
@@
$\min_{\widetilde{W}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ (55)
$+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}$
${\rm s.t.}{}{}$ $\left(V_{i2}^{\top}\otimes V_{i1}\right)\widetilde{W}=0,~{}~{}i=1,\dots,M.$
Recall
@@
$\min_{\widetilde{W}}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ (P)
$+\delta_{\Omega}(\widetilde{W})+\gamma\|(V_{2}\otimes V_{1})\widetilde{W}\|_{0}.$
@@
$\lim\limits_{\sigma\downarrow 0^{+}}f_{\sigma}(|y|)=I(y)=\left\{\begin{aligned} &0,~{}~{}y=0,\\ &1,~{}~{}\text{otherwise}.\end{aligned}\right.$
@@
$\min_{\widetilde{W}}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})$ ($\textbf{P}_{\sigma}$)
$+\delta_{\Omega}(\widetilde{W})+\gamma f_{\sigma}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}\right|\right).$
@@
[…]
@@
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}~{}\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})+\cdots$ ($\textbf{P}^{\prime}_{\sigma}$)
$+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W})+\gamma y^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\gamma g_{\sigma}^{*}(-y).$
@@
$r_{1}(\widetilde{W})=\langle{\rm vec}(R),\widetilde{W}\rangle+\delta_{\Gamma_{+}^{p}}(\widetilde{W})+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{1})$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\dots+\delta_{\Gamma_{+}^{n}}(\widetilde{\Psi}_{M})+\delta_{\Omega}(\widetilde{W}),$
@@
$\min_{\widetilde{W},y\in\mathbb{R}^{mn}_{+}}~{}H_{\sigma}(\widetilde{W},y)=r_{1}(\widetilde{W})+r_{2,\sigma}(y)+h(\widetilde{W},y).$ ($\textbf{P}^{\prime\prime}_{\sigma}$)
It is important to highlight that $r_{1},r_{2}$ are closed and convex, but $h$ is nonconvex and nonsmooth. It seems that, compared to optimization problem ($\textbf{P}_{\sigma}$), ($\textbf{P}^{\prime\prime}_{\sigma}$) remains a nonconvex problem but introduces auxiliary variable $y$, which even makes optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) more challenging. Interestingly, it is evident that $H_{\sigma}(\widetilde{W},y)$ is closed convex w.r.t. $\widetilde{W}$ and $y$, respectively. This observation suggests that by employing block coordinate descent (BCD), ($\textbf{P}^{\prime\prime}_{\sigma}$) can be converted to successive convex optimization problems. Specifically, given $y_{k-1}$, update $\widetilde{W}_{k}$; then given $\widetilde{W}_{k}$, update $y_{k}$, i.e.,
$\widetilde{W}_{k}\in\underset{\widetilde{W}\in\mathbb{R}^{p^{2}}}{\operatorname*{argmin}}~{}r_{1}(\widetilde{W})+h(\widetilde{W},y_{k-1}),$ ($\textbf{BP}\sigma-1$)
$y_{k}\in\underset{y\in\mathbb{R}_{+}^{mn}}{\operatorname*{argmin}}~{}r_{2,\sigma}(y)+h(\widetilde{W}_{k},y).$ ($\textbf{BP}\sigma-2$)
It is important to underline that both ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) may be classified as convex optimization problems. Our objective is then to solve ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternatively in order to approximate the global minimizers of nonconvex problem ($\textbf{P}^{\prime\prime}_{\sigma}$). Before discussing the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and ($\textbf{BP}\sigma-1$) ($\textbf{BP}\sigma-2$), we first propose the relation between ($\textbf{P}^{\prime\prime}_{\sigma}$) and (P). Denote
@@
Proof: Since for $x=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$
@@
[…]
Hence, theoretically, one can obtain the minimizers of optimization ($\textbf{P}^{\prime\prime}_{\sigma}$) by successively solving optimization problem ($\textbf{P}_{\sigma}$) with $\sigma\downarrow 0$. As mentioned above, we hope to solve problem ($\textbf{P}^{\prime\prime}_{\sigma}$) by solving ($\textbf{BP}\sigma-1$) and ($\textbf{BP}\sigma-2$) alternately, known as BCD method. However, due to the nonconvex and nonsmooth term $h(\widetilde{W},y)$, direct BCD may fail to converge and cause stable cyclic behavior [58]. It is worth to note that the BCD method can be regarded as a special case of BSUM framework (Algorithm 1); and hence, instead of BCD method, we will design suitable algorithm for solving ($\textbf{P}^{\prime\prime}_{\sigma}$) under a more general framework, i.e., BSUM framework.
The auxiliary function (see Definition 2) $u_{2,\sigma}(y,X_{k})$ can be selected as
$u_{2,\sigma}(y,X_{k})=\langle{\rm vec}(R),\widetilde{W}_{k}\rangle+r_{2,\sigma}(y)+h(\widetilde{W}_{k},y)$ (58)
with $X_{k}=(\widetilde{W}_{k},y_{k})$. Then, the subproblem
$\underset{y\in\mathbb{R}^{mn}_{+}}{\operatorname*{argmin}}~{}u_{2,\sigma}(y,X_{k})$
within BSUM framework is equivalent to problem ($\textbf{BP}\sigma-2$), which is further equivalent to the following problem
@@
Proof: Optimality condition implies that
@@
[…]
We then discuss the selection of $u_{1,\sigma}(\widetilde{W},X_{k})$. The proximal term is introduced, i.e.,
$u_{1,\sigma}(\widetilde{W},X_{k};\lambda)=$ $\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$ (61)
$+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2},$